OpenStudy (anonymous):

Write the function whose graph is the graph of y=x^3+3, but is reflected about the x-axis.

OpenStudy (anonymous):

Idk, but heres how to find it for the Y axis, try to reverse it to find the X axis: To find the eq of a function reflected over the y axis, replace x with -x in the function y= x^3 +3 y=(-x)^3 +3 y= -x^3 +3 If the original and reflected functions are the same, the function is symmetrical over the y axis.